OpenStudy (anonymous):

Pls Help me find the general solution of (1+x) (dy/dx) - xy = x+ x^2

OpenStudy (anonymous):

e^(x-ln|1+x|) y = integral of (e^x-ln|1+x|)) x

OpenStudy (anonymous):

Then I'm stuck with the integral of something I don't quite see

ganeshie8 (ganeshie8):

your integrating factor is correct and the left hand side is also correct

ganeshie8 (ganeshie8):

\[\large \int e^{x - \ln |1+x|}x ~dx\]

ganeshie8 (ganeshie8):

are you stuck with that integral ?

OpenStudy (anonymous):

Yes!

ganeshie8 (ganeshie8):

use exponent laws

ganeshie8 (ganeshie8):

\[\large a^{m-n} = \dfrac{a^m}{a^n}\] etc

OpenStudy (anonymous):

Then I just do integration by parts?

ganeshie8 (ganeshie8):

\[\large \int e^{x - \ln |1+x|}x ~dx = \int \dfrac{e^x}{e^{\ln |1+x|}}x ~dx = \int \dfrac{xe^x}{1+x} ~dx\]

ganeshie8 (ganeshie8):

yeah, maybe we can simplify a bit before IBP : \[\large = \int \dfrac{(1+x-1)e^x}{1+x}~dx\]

ganeshie8 (ganeshie8):

\[\large = \int e^x - \dfrac{e^x}{1+x}~dx\]

ganeshie8 (ganeshie8):

not possible in elementary functions http://www.wolframalpha.com/input/?i=%5Cint+e%5Ex%2F%281%2Bx%29+dx so you're lucky, notthing to do here, just leave the integral as it is

OpenStudy (anonymous):

Oh wow, thanks

ganeshie8 (ganeshie8):

wait a sec, your integrating factor doesn't look correct :o

ganeshie8 (ganeshie8):

\[\large (1+x) \dfrac{dy}{dx} - xy = x+ x^2\]

ganeshie8 (ganeshie8):

\[\large \dfrac{dy}{dx} + \dfrac{ -x}{1+x}y = x\]

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats what i have

ganeshie8 (ganeshie8):

IF = \[\large e^{\int \dfrac{-x}{1+x} ~dx} = e^{-\int \dfrac{1+x-1}{1+x} ~dx} \]

ganeshie8 (ganeshie8):

\[ \large e^{-\int 1 - \dfrac{1}{1+x} ~dx} = e^{-x + \ln |1+x|}\]

ganeshie8 (ganeshie8):

right ?

ganeshie8 (ganeshie8):

Notice that your signs were flipped in the exponent thats the reason we arrived at a messedup integral

OpenStudy (anonymous):

Oh I didnt carry the - sign

ganeshie8 (ganeshie8):

simplify the IF before multiplying both sides : \[ \large e^{-\int 1 - \dfrac{1}{1+x} ~dx} = e^{-x + \ln |1+x|} = \dfrac{1+x}{e^x}\]

ganeshie8 (ganeshie8):

you may be able to integrate easily with this IF :) good luck !

OpenStudy (anonymous):

Thanks so much!!

ganeshie8 (ganeshie8):

yeah wolfram says its integrable http://www.wolframalpha.com/input/?i=%5Cint+x%281%2Bx%29%2Fe%5Ex+dx